OpenStudy (anonymous):

test question need help Violet was carrying around the steps to solve Square root of 3x plus 1 = 4. On her way through the Math Factory, Violet dropped them and they mixed with other steps. Put the correct steps in order to solve Square root of 3x plus 1 = 4.

OpenStudy (anonymous):

A. Subtract 1 from both sides B. Subtract 4 from both sides C. Square root both sides D. Square both sides E. Divide by 3 from both sides F. Divide by 4 from both sides

OpenStudy (anonymous):

A.)First A, then E, then D B.)First D, then B, then A C.)First D, then A, then E D.) First C, then E, then A

OpenStudy (ddcamp):

\[\sqrt{3x}+1\ or\ \sqrt{3x+1}\]

OpenStudy (anonymous):

its \[\sqrt{3x+1}\]

OpenStudy (ddcamp):

OK. We can't work with the variable x while it's under the square root sign. What would we do to get rid of the square root sign?

OpenStudy (anonymous):

umm multiply it on both sides by 3? (im clueless on this btw)

OpenStudy (ddcamp):

We get rid of square root signs by squaring both sides: \[\sqrt{3x+1}^2 = 4^2 \\ \rightarrow 3x+1 = 16 \]

OpenStudy (anonymous):

then we find x?

OpenStudy (ddcamp):

Yes

OpenStudy (anonymous):

so subtract 1from both sides and divide by 3?

OpenStudy (anonymous):

Square both sides, subtract 1 from both sides, and divide by 3 from both sides.